The width of a rectangle is 2 meters larger than its height. The diagonal brace measures sqrt 6m. Find the width and height.
Answer by stanbon(48568)

The width of a rectangle is 2 meters larger than its height. The diagonal brace measures sqrt 6m. Find the width and height.
----------------
Let height be "x".
Then width is "x+2"
----------
Pythagoras says x^2 + (x+2)^2 = (sqrt(6))^2
x^2 + x^2 + 4x + 4 = 6
2x^2 + 4x -2 = 0
x^2 + 2x -1 = 0
x = [-2 +- sqrt(4 - 4*1*-1)]/2
x = [-2 +- sqrt(8)]/2
x = [-1 +- sqrt(2)]
Positive solution:
x = -1 + 1.414 = 0.414 meters (height)
x+2 = 2.414 meters (width)
===========================
Cheers,
Stan H.